Suppose you heat 1,000 g of water to boiling and mix it with 3,000 g of water at 160C contained in an aluminum pail of mass 50 g. What is the temperature of the resulting mixture? Show your complete solution.



1
Expert's answer
2021-11-21T17:25:21-0500

"cm_1(t-t_0)=cm_2(t_0-16\u00b0)+c_Am_A(t_0-16\u00b0)"


"4200\cdot 1\cdot(100-t_0)=4200\cdot 3\cdot(t_0-16°)+920\cdot 0.05\cdot(t_0-16°)\to"


"t_0\approx37°" . Answer